# Vertex-Minimal Paper Tori

### v4

@@ 5.4 The Final Estimates
For each relevant pair $(V_{1},V_{2})$ of normalized vectors (in the notation above) we have $\|V_{1}\|,\|V_{2}\|\in[0.8,2.4]$ and $|u_{2}|>0.016$.
Proof: The norm bounds combine a direct calculation for $T$ with the observation that the norms change by at most $10^{-28}$ when we move to some other point in $B$.
The bound on $u_{2}$ is more subtle. For each “raw pair” $(V_{1}^{\prime},V_{2}^{\prime})$ associated to our triangle we compute
$a(V_{1}^{\prime},V_{2}^{\prime})=|u_{1}^{\prime}v_{2}^{\prime}-u_{2}^{\prime}v_{1}^{\prime}|.$
This quantity measures the area of the parallelogram spanned by our vectors when we project them into the plane. In particular $a(V_{1}^{\prime},V_{2}^{\prime})=a(V_{1},V_{2})$ when $(V_{1},V_{2})$ is the normalized pair associated to $(V_{1}^{\prime},V_{2}^{\prime})$. We check directly that $|a(V_{1}^{\prime},V_{2}^{\prime})|>0.02$ for all relevant triangles. We also see from Equation 1 that the projection of each vector of $T$ into the plane has length less than $1.2$. Hence $|v_{1}|<1.2$. Combining everything, we have
$|u_{2}|=\frac{|u_{2}||v_{1}|}{|v_{1}|}=\frac{|a(V_{1},V_{2})|}{|v_{1}|}=\frac{|a(V_{1}^{\prime},V_{2}^{\prime})|}{|v_{1}|}>\frac{0.02}{1.2}>0.016.$
This completes the proof. $\spadesuit$
To make our proof more robust, we will work with worse bounds:
$\|V_{1}\|^{2},\|V_{2}\|^{2}\in(10^{-1},10)\hskip 30.0pt|u_{2}|>10^{-2}.$ (20)
Case 1: We have
$g_{12}=\frac{u_{2}^{4}v_{1}^{4}}{\left(u_{2}^{2}(v_{1}^{2}+w_{1}^{2})+(v_{2}w_{1}-v_{1}w_{2})^{2}\right)^{3}}<\frac{10^{4}}{(10^{-4}\times 10^{-2})^{3}}<10^{46}.$ (21)
Case 2: We have
$g_{11}=\frac{v_{1}^{2}\left(-2w_{1}(v_{2}w_{1}-v_{1}w_{2})^{3}+u_{2}^{2}(v_{1}^{2}+w_{1}^{2})(v_{1}^{2}v_{2}-2v_{2}w_{1}^{2}+3v_{1}w_{1}w_{2})\right)^{2}}{(v_{1}^{2}+w_{1}^{2})^{4}\left(u_{2}^{2}(v_{1}^{2}+w_{1}^{2})+(v_{2}w_{1}-v_{1}w_{2})^{2}\right)^{3}}<^{*}$
$\frac{32\times 80\times 10^{16}}{(10^{-1})^{4}\times(10^{-4}\times 10^{-1})^{3}}<\frac{10^{20}}{10^{-19}}<10^{46}.$ (22)
The starred inequality comes from the fact that when we expand the numerator out completely it has $32$ terms, all coefficients less than or equal to $80$ in absolute value, and maximum degree $16$.
Case 3: We have
$g_{22}=\frac{\left(\begin{aligned} &2v_{2}w_{2}(v_{2}w_{1}-v_{1}w_{2})^{3}+u_{2}^{4}(v_{1}^{3}v_{2}+3v_{1}^{2}w_{1}w_{2}+2w_{1}^{3}w_{2})\\ &+u_{2}^{2}\big{(}4v_{2}^{2}w_{1}^{3}w_{2}-6v_{1}v_{2}w_{1}^{2}w_{2}^{2}+3v_{1}^{2}w_{1}w_{2}(v_{2}^{2}+w_{2}^{2})+v_{1}^{3}(v_{2}^{3}-v_{2}w_{2}^{2})\big{)}\end{aligned}\right)^{2}}{\left(u_{2}^{2}+v_{2}^{2}+w_{2}^{2}\right)^{4}\left(u_{2}^{2}(v_{1}^{2}+w_{1}^{2})+(v_{2}w_{1}-v_{1}w_{2})^{2}\right)^{3}}<^{*}$
$\frac{55\times 124\times 10^{16}}{(10^{-1})^{4}\times(10^{-4}\times 10^{-1})^{3}}<\frac{10^{20}}{10^{-19}}<10^{46}.$ (23)
The starred inequality has the same explanation as in Case 2.
@@
6 Appendix: Coordinates and Pictures
6.1 Another Pup Tent
After writing the first version of this paper, I fooled around with the coordinates and found coordinates for a pup tent that are simpler than the ones in Equation 1. Here are the coordinates up to $32$ digits.
[…]
The way I found this example is that I had ChatGPT-5 rotate the original example around the $Z$-axis, looking for rotations which can be truncated to two decimal places with only small changes. So, for instance, perhaps the first coordinate was .7803.7803, and there was not much to change. I certainly could have done this without the help of ChatGPT, but I was feeling lazy.
The reason to have very simple coordinates, aside from the aesthetic value, would be to try to recognize the $z_{0},z_{1},z_{2}$ coordinates as algebraic numbers. This would be nice, because then one could have an explicit $8$-vertex paper torus rather than one defined by the Implicit Function Theorem. I have not had any success doing this, however.
6.2 Two More Pictures
The next picture we show is a layout for building your own pup tent. The figure gives a folding guide and the fold lines for a fundamental domain of the torus. The picture is decorated with one of the level set foliations. When the thing is correctly folded and suitably rotated, the level curves all lie in horizontal planes.
The final picture shows some $3D$ pictures that Peter Doyle made. The left and middle columns show two annuli that together form the torus in the right column. The rows correspond to viewing the tent from the top, bottom, broadside, and head on.

### v9

@@ 3.3 Proof of the Crude Bound Lemma
For each relevant pair $(V_{1},V_{2})$ of vectors,
$\|V_{1}\times V_{2}\|^{2},\|V_{1}\|^{2},\|V_{2}\|^{2}\in(10^{-1},10).$
Proof: For $T$ we get the bounds $\|V_{k}\|\in[0.48,2.4]$ and $\|V_{1}\times V_{2}\|\in[.85,2.4]$. These various quantities hardly change at all when we switch to a different relevant torus and change our coordinates by less than $10^{-4}$. $\spadesuit$
Case 1: Let $W_{k}=(u_{k},v_{k},0)$ be the projection of $V_{k}$ to the $XY$-plane. We have
$g_{21}=g_{12}=\frac{\|W_{1}\times W_{2}\|^{4}}{\|V_{1}\times V_{2}\|^{6}}\leq\frac{\|V_{1}\times V_{2}\|^{4}}{\|V_{1}\times V_{2}\|^{6}}=\frac{1}{\|V_{1}\times V_{2}\|^{2}}<10.$
The second inequality comes from Lemma 3.7.
Case 2: With $W_{k}$ defined as above, we have
$g_{11}=\frac{\bigg(t_{1}w_{1}w_{2}+t_{2}w_{1}w_{2}^{3}+t_{3}w_{1}^{3}w_{2}+t_{4}w_{1}^{2}w_{2}^{2}+t_{5}w_{1}^{4}+t_{6}w_{1}^{2}+t_{7}\bigg)^{2}}{\|V_{1}\|^{8}\|V_{1}\times V_{2}\|^{6}},$
$t_{1}$ $=3\,\|W_{1}\|^{2}\,\|W_{1}\times W_{2}\|^{2},$ $\quad t_{2}$ $=2\,\|W_{1}\|^{4},$
$t_{3}$ $=3\,(W_{1}\cdot W_{2})^{2}\;+3\,\|W_{1}\|^{2}\|W_{2}\|^{2},$ $\quad t_{4}$ $=-6\,\|W_{1}\|^{2}(W_{1}\cdot W_{2}),$
$t_{5}$ $=-2\,\|W_{2}\|^{2}(W_{1}\cdot W_{2}),$ $\quad t_{6}$ $=-\,\|W_{1}\times W_{2}\|^{2}\,(W_{1}\cdot W_{2}),$
$t_{7}$ $=\|W_{1}\|^{2}\,\|W_{1}\times W_{2}\|^{2}\,(W_{1}\cdot W_{2}).$
Write this as $g_{11}=N^{2}/D$. Using Lemma 3.7 and the fact that $|w_{1}|,|w_{2}|<1$ we see that $|N|<10^{7/2}$ and $D>10^{-7}$. Hence $g_{11}<10^{14}$. A similar calculation works for $g_{22}$ and indeed this case follows from that of $g_{11}$ and symmetry.
@@
4 Discussion and Pictures
4.1 Experimentation
Recall that there are $7$ combinatorial types of triangulation of the torus having $8$ vertices. Figure 1.2 shows the best of these, the one with uniform degree. For all but the worst triangulation, I generated lots of embedded examples by randomly sampling collections of $8$ points on the unit sphere and testing the embedding condition.
Next, I used a hill climbing algorithm. The way this works is that you want to minimize an objective function on a space. You start at a random point and then evaluate the function. You then make a random perturbation and move to the new location if the function is lower. If you are in a smooth manifold and the function is smooth and nonsingular, then the random perturbations often move in the direction of the gradient. In this setting the hill climbing algorithm somewhat mimics gradient flow. This approach works pretty well in low dimensions.
My hill-climbing algorithm started with random embedded examples on the unit sphere and repeatedly perturbed them – not necessarily keeping them on the unit sphere – with the goal of minimizing an objective flatness function. The flatness function computes the maximum deviation of a cone angle from being $2\pi$.
After a lot of experimentation, I implemented $4$ tricks that helped immensely.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    It helped to force the examples to have $2$-fold rotational symmetry. This simplification came later. After I found some asymmetric examples, Samuel Lelièvre asked me about symmetric examples and I tweaked the program. The program works more reliably after incorporating the symmetry.
4.2 Pictorial Study
With all the tricks above in place, and a lot of playing around, I found some embedded and $2$-fold symmetric examples of face number $6$ which are flat up to $10^{-16}$. I took the nicest one and simplified the coordinates, feeding it back into the program to make it near-flat again. After some fooling around with the digits I arrived at the example in Equation 1.
Figure 4.1 shows three Mathematica plots of the pup tent, from various angles.
Figure 4.1: 3D plots of the pup tent
Figure 4.2 shows a numerical computation of the intrinsic geometric structure of the faces of the pup tent. We are showing part of the universal cover of the triangulation. We have also highlighted the fundamental domain that corresponds to the one in Figure 1.2. The blue triangles correspond to the faces of the pup tent on the convex hull boundary. There are $6$ of them in the fundamental domain.
Figure 4.2: The intrinsic structure.
Figure 4.3 shows a second coloring of the same fundamental domain. We have labeled the points. We list out numerical approximations to the $8$ blue triangles. We get the pink ones by reflecting the blue ones in the origin and applying the permutation $j\to 7-j$ to the labels.
Figure 4.3: A fundamental domain for the tiling
$\begin{array}[]{lll}\{6,0,1\}\!:&\{-2.9340654093-0.6455398803i,\;-1.8673131665-0.4088246583i,\;-2.3837468425+1.5965906856i\},\\[2.0pt] \{3,1,0\}\!:&\{-0.8925492572+0.4429076765i,\;-2.3837468425+1.5965906856i,\;-1.8673131665-0.4088246583i\},\\[2.0pt] \{3,4,1\}\!:&\{-0.8925492572+0.4429076765i,\;-1.6100587412+2.3020419213i,\;-2.3837468425+1.5965906856i\},\\[2.0pt] \{2,0,7\}\!:&\{-0.3555738883-0.4429076765i,\;-1.8673131665-0.4088246583i,\;-0.9478910867-1.3850741341i\},\\[2.0pt] \{3,0,2\}\!:&\{-0.8925492572+0.4429076765i,\;-1.8673131665-0.4088246583i,\;-0.3555738883-0.4429076765i\},\\[2.0pt] \{3,2,5\}\!:&\{-0.8925492572+0.4429076765i,\;-0.3555738883-0.4429076765i,\;+0.3555738883+0.4429076765i\},\\[2.0pt] \{3,5,6\}\!:&\{-0.8925492572+0.4429076765i,\;+0.3555738883+0.4429076765i,\;-0.1188611560+1.1483589122i\},\\[2.0pt] \{5,0,6\}\!:&\{+0.3555738883+0.4429076765i\;+1.8673131665+0.40882465835i,\;+0.9478910867+1.3850741341i\}\end{array}$
Figures 4.2 and 4.3 look quite tame, but the extrinsic geometry of the pup tent is beautiful and very intricate. Figure 4.4 shows one of the slices of the pup tent by a plane parallel to the $Z$ axis. The solid handlebody bounded by the torus is colored orange. This slice is very spiky and sort of “barely embedded”, but the embeddedness is robust compared to the plot precision.
Figure 4.4: A horizontal slice of the pup tent
Figure 4.5 shows the slice by the $XZ$ plane, which goes through the axis of symmetry of the torus. The slice of the handlebody is an annulus. We have included a closeup of the bottom so that the reader can see it better.
Figure 4.5: The $XZ$ planar slice of the pup tent.
Figure 4.6 shows the slice by the $YZ$ plane, with a closeup of the bottom.
Figure 4.6: The $YZ$ planar slice of the pup tent
The left side of Figure 4.7 shows a random slice of the pup tent.
Figure 4.7: Extrinsic and Intrinsic structures compared
My supplementary notes [S] give additional information about the pup tent, including a pattern than you can cut out and (with effort) fold into one. Also, my Java program lets you see many more pictures like this.